article "Standard Costing Is Alive and Well at Parker Brass" by D. Johnsen and P. Sopariwala, Management Accounting Quarterly (Winter 2000), pp. 12-20.

PANEL B: WORKINGS

Setups

Production

Total

Standard time per unit:

Standard setup time ( hours)

4

Standard production time (hours)

40

Standard batch size (units)

2,000

2,000

Hence, standard time per unit (hours)

0.002

0.020

0.022

(a) Given Information as above, if # of units actually produced is 1,000 units, compute actual setup hours for 1 setup, actual productive labor hours to make 1,000 units, and actual labor cost for 1,000 units at $10 per labor hour. Please show your computation steps.

Actual quantity produced (units)

1,000

Actual setup hours for 1 setup

?

Actual productive labor hours to make 1,000 units

?

Actual labor cost for 1,000 units at $10 per labor hour

$?

(b) Given Information as Panel A and Panel B above, if # of units actually produced is 1,000 units, compute standard time charged (hours) for Setups and compute standard time charged (hours) for Production. Please show your computation steps.

Setups

Production

Total

Standard time charged for 1,000 units:

# of units actually produced

1,000

1,000

1,000

Standard time charged (hours)

?

?

?
Answer #1

a. Standard setup time = 4 hours (given in Panel B)

Standard batch size = 2000 units (given in Panel B)

Actual quantity produced (units) = 1000 (given)

Actual setup hours for 1 setup = Standard setup time * (Actual quantity produced / Standard batch size)

= 4 * (1000 / 2000) = 2 hours

Standard production time = 40 hours (given in Panel B)

Standard batch size = 2000 units (given in Panel B)

Actual quantity produced (units) = 1000 (given)

Actual productive hours for 1000 units = Standard production time * (Actual quantity produced / Standard batch size)

= 40 * (1000 / 2000) = 20 hours

Actual labor cost for 1,000 units at $10 per labor hour = Actual productive hours for 1000 units * Rate of labor

= 20 hours * $10 = $ 200

b.

Setups

Production

Total

Standard time per unit (as given in Panel B):

0.002 0.02 0.022

# of units actually produced

1,000

1,000

1,000

Standard time charged (hours) = Standard time per unit * Units actually produced

2

20

22